Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 19:12:18 on February 22, 2009
Entry number 262711
This entry is a followup to: 262534
Followups:
keyword=Experiment Summary for Feb 21, Day - Feb 22, Day
Summary for Feb 21 day - Feb 22 day
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/02/21 day | 6.33 |
0.51 | 1.17 |
2009/02/21 swing | 7.23 |
0.20 | 0.57 |
2009/02/22 owl | 7.34 |
0.14 | 0.53 |
2009/02/22 day | 4.11 |
0.12 | 3.75 |
total | 25.0 |
1.0 | 6.0 |
Main events/issues:
- February 21st
- Transverse production
- 14:40 Completed kinematic H: Plhrs = 1.27 GeV/c,
45o.
- 15:00 Began kinematic F: Plhrs = 1.13 GeV/c,
45o.
- Smooth Production the rest of the day.
- Average target polarization: ~ 61.8% from day shift 2/21
through swing shift 2/21.
- February 22nd
- 10:45 Completed kinematic F: 0.85 C out of 0.86 C.
- 10:55 Start kinematic D: Plhrs = 0.9 GeV/c,
45o.
- 12:49-15:37 No beam due to injector and North LINAC dropped
out of HV
then after restoration no beam due to 1L16 HPA
amplifier fault.
- 15:37-16:00 Beam restored and beam profile checks begin.
- Average target polarization: ~ 61.7% from owl shift 2/22
through day shift 2/22.
Plans for the next day:
- Production for kinematic D.
- At 18:00 target spin rotation to longitudinal.
- Longitudinal production running during kinematic D on left HRS.
- February 23rd from 09:00 to 13:00
- RF recovery
- Controlled access to investigate BigBite shower hardware
- Pass change to fourth
- Target spin rotation to transverse
- Begin unpolarized production at 4th pass for left HRS.
A copy of this log entry has been emailed to: brads,jpchen,jiang,meziani,choi